# A SAT-centered XAI method for Deep Learning based Video Understanding

### v1

@@ 5.1.2 Addressing the Limitations of Current Deep Learning Models in Terms of Explainability
The inherent complexity of deep learning models often makes it difficult to understand their decision-making processes [48, 72]. The proposed method attempts to overcome this by providing a different lens through which to analyze these models – the lens of formal logic [73, 39]. By successfully encoding a deep learning model into a SAT problem, we are essentially creating a symbolic representation of its behavior [46, 40]. The explanations generated by querying this symbolic representation using a SAT solver can offer insights into the model’s logic that might be hard to obtain by directly analyzing the model’s weights and activations [45].
For instance, the minimal set of necessary input features identified for a prediction can highlight which aspects of the video the model deemed crucial [76]. The minimal changes required to flip a prediction can reveal the model’s sensitivity to certain features and the decision boundaries it has learned [17, 64]. This level of insight can be valuable for understanding the model’s strengths and weaknesses, identifying potential biases, and guiding efforts to improve its performance or robustness [67, 66].
@@ 5.1.3 Potential Applications and Impact of the Proposed Model
The proposed explainable video understanding model has the potential to be applied in various domains where understanding the reasoning behind a model’s predictions is critical [4]. In surveillance, for example, it could explain why a particular activity was flagged as suspicious, potentially highlighting the specific movements or interactions that led to the alert [7, 11]. In autonomous driving, it could provide explanations for why the system made a certain driving decision in a complex scenario, which could be crucial for safety and accountability [31, 43]. In medical diagnosis from video data (e.g., analyzing patient movements or surgical procedures), the model could explain the reasoning behind a diagnosis, aiding medical professionals in understanding and trusting the AI’s assessment [3, 5]. In content analysis, it could explain why a video was classified in a certain way, providing insights into the features the model considers important for different categories [9, 56].
The ability to provide formal explanations can significantly impact the trust and accountability associated with using deep learning in these areas [38, 19]. When users can understand the logical reasons behind a model’s predictions, they are more likely to trust its decisions [63, 30]. Furthermore, if errors occur, the explanations can help in debugging the model and identifying the root causes of the failures, leading to more effective improvements [61, 21].
@@ 5.1.4 Challenges and Limitations of the Proposed Approach
Despite its potential advantages, the proposed method also faces several challenges and limitations that need to be considered [28, 29].
@@
Encoding a deep learning model, especially a large and complex one used for video understanding, into a SAT problem can be computationally very expensive [35, 32]. The number of Boolean variables and clauses required to represent the model’s weights, biases, activations, and operations can grow rapidly with the size and complexity of the network [36, 24]. Similarly, encoding a video, even after feature extraction, can also lead to a large number of variables and clauses, especially when considering temporal dependencies [20, 37].
Solving the resulting SAT problem can also be computationally intensive, as SAT is an NP-complete problem [2, 69]. While modern SAT solvers have made significant progress and can handle instances with millions of variables and clauses, the size and complexity of the SAT problems arising from encoding deep learning models for video might still pose a significant scalability challenge [34]. Future work needs to explore efficient encoding techniques and investigate the use of optimized SAT solvers, potentially including parallel solvers, to address these issues. Abstractions or approximations in the encoding process might also be necessary to achieve practical scalability.
@@ 6.1 Directions for Future Work
Several avenues exist for future research to build upon and extend the proposed method [78]:
  • Exploring Different SAT Solving Techniques and Optimizations: Investigating the effectiveness of various SAT solving algorithms, including CDCL with different heuristics, and exploring optimizations tailored to the specific types of SAT problems generated from encoding deep learning models for video [79, 70].
  • Extending the Model to Handle Different Types of Video Understanding Tasks: Applying and adapting the proposed method to a wider range of video understanding tasks beyond action recognition, such as video captioning, anomaly detection, and event recognition. Each task might require specific considerations in terms of feature representation and explanation queries [47, 13].
  • Investigating Methods for Generating More Human-Interpretable Explanations from the SAT-Based Framework: Developing techniques to translate the logical explanations obtained from the SAT solver into more human-readable formats, such as natural language descriptions, visual summaries, or high-level logical rules. This would make the explanations more accessible and useful to end-users [54, 71, 14].
  • Potential for Empirical Evaluation in Specific Video Understanding Domains: Conducting empirical evaluations of the proposed model on real-world video datasets to assess its effectiveness in providing meaningful explanations and to evaluate its scalability for practical applications. This would involve selecting appropriate deep learning models for video understanding, designing relevant explanation queries, and analyzing the quality and computational cost of the generated explanations [15, 41, 27].
  • Exploring Alternative Logical Frameworks: Investigating the use of more expressive logical frameworks like SMT or first-order logic to encode deep learning models for video, potentially allowing for a more direct representation of continuous values and complex relationships. This would involve addressing the challenges associated with solving problems in these more expressive logics [62, 65, 48].
@@ 6.2 Conclusion
This paper has presented a novel formal SAT-based explanation model for deep learning in video understanding [72, 73]. The proposed method aims to address the limitations of existing XAI techniques by integrating the rigorous reasoning capabilities of SAT solvers with the principles of formal explainable AI [39, 46, 40, 45]. By formally encoding deep learning models and video data into a logical framework and formulating explanation queries as satisfiability problems, the method offers the potential to generate logic-based explanations with formal guarantees [76, 17, 64, 67, 66]. The paper detailed the conceptual framework, the process of encoding deep learning models and video data, the formulation of ”Why?” and ”Why not?” questions, and a novel architecture integrating a SAT solver with a deep learning video understanding model [4, 7, 11, 31, 43]. While challenges related to computational complexity and the representational power of propositional logic remain, the proposed approach offers a promising direction for enhancing the explainability of deep learning in the complex and critical domain of video understanding [3, 5, 9, 56, 38, 19]. Future work will focus on addressing these challenges, extending the model to various video understanding tasks, and developing methods for generating more human-interpretable explanations, ultimately aiming to increase trust and accountability in AI-driven video analysis [63, 30, 61, 21, 28, 29].

### v2

@@ 5.1.2 Addressing the Limitations of Current Deep Learning Models in Terms of Explainability
The inherent complexity of deep learning models often makes it difficult to understand their decision-making processes. The proposed method attempts to overcome this by providing a different lens through which to analyze these models – the lens of formal logic. By successfully encoding a deep learning model into a SAT problem, we are essentially creating a symbolic representation of its behavior. The explanations generated by querying this symbolic representation using a SAT solver can offer insights into the model’s logic that might be hard to obtain by directly analyzing the model’s weights and activations.
For instance, the minimal set of necessary input features identified for a prediction can highlight which aspects of the video the model deemed crucial. The minimal changes required to flip a prediction can reveal the model’s sensitivity to certain features and the decision boundaries it has learned. This level of insight can be valuable for understanding the model’s strengths and weaknesses, identifying potential biases, and guiding efforts to improve its performance or robustness.
@@ 5.1.3 Potential Applications and Impact of the Proposed Model
The proposed explainable video understanding model has the potential to be applied in various domains where understanding the reasoning behind a model’s predictions is critical. In surveillance, for example, it could explain why a particular activity was flagged as suspicious, potentially highlighting the specific movements or interactions that led to the alert. In autonomous driving, it could provide explanations for why the system made a certain driving decision in a complex scenario, which could be crucial for safety and accountability. In medical diagnosis from video data (e.g., analyzing patient movements or surgical procedures), the model could explain the reasoning behind a diagnosis, aiding medical professionals in understanding and trusting the AI’s assessment. In content analysis, it could explain why a video was classified in a certain way, providing insights into the features the model considers important for different categories.
The ability to provide formal explanations can significantly impact the trust and accountability associated with using deep learning in these areas. When users can understand the logical reasons behind a model’s predictions, they are more likely to trust its decisions . Furthermore, if errors occur, the explanations can help in debugging the model and identifying the root causes of the failures, leading to more effective improvements.
@@ 5.1.4 Challenges and Limitations of the Proposed Approach
Despite its potential advantages, the proposed method also faces several challenges and limitations that need to be considered.
@@
Encoding a deep learning model, especially a large and complex one used for video understanding, into a SAT problem can be computationally very expensive. The number of Boolean variables and clauses required to represent the model’s weights, biases, activations, and operations can grow rapidly with the size and complexity of the network. Similarly, encoding a video, even after feature extraction, can also lead to a large number of variables and clauses, especially when considering temporal dependencies.
Solving the resulting SAT problem can also be computationally intensive, as SAT is an NP-complete problem. While modern SAT solvers have made significant progress and can handle instances with millions of variables and clauses, the size and complexity of the SAT problems arising from encoding deep learning models for video might still pose a significant scalability challenge. Future work needs to explore efficient encoding techniques and investigate the use of optimized SAT solvers, potentially including parallel solvers, to address these issues. Abstractions or approximations in the encoding process might also be necessary to achieve practical scalability.
@@ 6.1 Directions for Future Work
Several avenues exist for future research to build upon and extend the proposed method:
  • Exploring Different SAT Solving Techniques and Optimizations: Investigating the effectiveness of various SAT solving algorithms, including CDCL with different heuristics, and exploring optimizations tailored to the specific types of SAT problems generated from encoding deep learning models for video.
  • Extending the Model to Handle Different Types of Video Understanding Tasks: Applying and adapting the proposed method to a wider range of video understanding tasks beyond action recognition, such as video captioning, anomaly detection, and event recognition. Each task might require specific considerations in terms of feature representation and explanation queries.
  • Investigating Methods for Generating More Human-Interpretable Explanations from the SAT-Based Framework: Developing techniques to translate the logical explanations obtained from the SAT solver into more human-readable formats, such as natural language descriptions, visual summaries, or high-level logical rules. This would make the explanations more accessible and useful to end-users.
  • Potential for Empirical Evaluation in Specific Video Understanding Domains: Conducting empirical evaluations of the proposed model on real-world video datasets to assess its effectiveness in providing meaningful explanations and to evaluate its scalability for practical applications. This would involve selecting appropriate deep learning models for video understanding, designing relevant explanation queries, and analyzing the quality and computational cost of the generated explanations.
  • Exploring Alternative Logical Frameworks: Investigating the use of more expressive logical frameworks like SMT or first-order logic to encode deep learning models for video, potentially allowing for a more direct representation of continuous values and complex relationships. This would involve addressing the challenges associated with solving problems in these more expressive logics.
@@ 6.2 Conclusion
This paper has presented a novel formal SAT-based explanation model for deep learning in video understanding. The proposed method aims to address the limitations of existing XAI techniques by integrating the rigorous reasoning capabilities of SAT solvers with the principles of formal explainable AI. By formally encoding deep learning models and video data into a logical framework and formulating explanation queries as satisfiability problems, the method offers the potential to generate logic-based explanations with formal guarantees. The paper detailed the conceptual framework, the process of encoding deep learning models and video data, the formulation of ”Why?” and ”Why not?” questions, and a novel architecture integrating a SAT solver with a deep learning video understanding model. While challenges related to computational complexity and the representational power of propositional logic remain, the proposed approach offers a promising direction for enhancing the explainability of deep learning in the complex and critical domain of video understanding. Future work will focus on addressing these challenges, extending the model to various video understanding tasks, and developing methods for generating more human-interpretable explanations, ultimately aiming to increase trust and accountability in AI-driven video analysis.